OpenStudy (anonymous):

Henry grabs one shoe at a time at random without replacement from a box of shoes in his closet. The box contains 2 brown shoes, 4 black shoes, and 4 sneakers. What is the probability that the first two shoes he chooses will be brown? Round to the nearest percent if necessary.

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

2/10*1/9

OpenStudy (anonymous):

good catch

OpenStudy (kropot72):

\[P(2brown)=\frac {\left(\begin{matrix}2 \\ 2\end{matrix}\right)\left(\begin{matrix}8 \\ 0\end{matrix}\right)}{\left(\begin{matrix}10 \\ 2\end{matrix}\right)}\] \[P(2brown)=\frac{2}{10\times 9}\]

OpenStudy (kropot72):

@amberose10 Have you calculated the answer yet?

OpenStudy (campbell_st):

the easiet way to do the question is P(brown shoe) = 2/10 1st selection then there are only 9 shoes left and 1 brown show P(brown shoe) = 1/9 2nd selection so P( Brown, Brown) = 2/10 x 1/9

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

i need help with this question

OpenStudy (anonymous):

ill give a medal